# Comparing the Robustness of Simple Network Scale-Up Method (NSUM) Estimators

### 08 Dec 2023 v2

@@ 1 Introduction
Surveys are a standard approach to estimating the size of subpopulations, or groups of people with a particular trait. Many key subpopulations of interest, however, are hard to reach***We note that researchers and community partners have raised concerns about the term “hard-to-reach”; we use this term here to encompass communities that match one or more of the above conditions, and we echo the message of routen2022HTR that the onus of reaching such populations is on researchers and funders, not on the populations themselves. with standard surveys for a number of possible reasons: (1) it may be hard to get a list of the members of this subpopulation or they may be hard to contact, as in the case of the homeless subpopulation; (2) it may be hard to accurately determine their membership in the subpopulation because their group status is stigmatized, as in the case of heavy drug users; and/or (3) the subpopulation may be rare in the general population (Bernard et al., 1991; Killworth et al., 1998). The limitations of standard survey methodology in reaching these groups has motivated the development of better adapted methods.
@@
The other strategy, the network scale-up method (NSUM), is to conduct a traditional survey with a representative sample from the general or frame population and ask respondents how many people they know in the hard-to-reach populations of interest, then use information about their personal network sizes (degrees) to scale up their data into an estimate for the general prevalence of those populations (Bernard et al., 1991, 2010). Researchers could ask respondents how many people they know (their degree) and how many people they know in the hard-to-reach group. However, people often cannot provide an accurate estimate for their degree (Killworth2003; McCormick et al., 2010), leading to inaccurate hard-to-reach population prevalence estimates. Therefore, respondents’ degrees are often estimated by asking other questions that are used to estimate the degree. In the probe group method, respondents are asked how many people they know in various groups of known size that collectively are thought to be representative of the general population (Killworth et al., 1998; McCarty2001comparing; McCormick et al., 2010). The responses to questions of the form “How many people do you know with X trait?” are known as aggregated relational data or ARD (McCormick and Zheng, 2015). This approach does not require knowing whether the respondents themselves are in the hard-to-reach populations.
In this paper we focus on the NSUM strategy. A recent literature review by Ocagli2021 suggests the majority of studies use one of the simplest estimators. However, we find that when a key model assumption is violated by the presence of barrier effects, which may often occur in practice, the mean squared error (MSE) may actually tend to be much smaller for another simple estimator that is equally easy to implement but much less commonly used. We demonstrate this through theoretical derivations as well as simulated surveys on a real network data set.
@@ 2 NSUM estimators
The target of estimation is the prevalence $r$ of the hard-to-reach group $H$ in the general population, which equals the ratio of the group size $N_{H}$ to the size $N$ of the general or frame population. For each respondent in a sample of size $n$, let $Y_{iH}$ represent respondent $i$’s response to the ARD question, “How many people do you know in $H$?” Let $D_{i}$ represent the degree of respondent $i$. In principle, we could determine $D_{i}$ by directly asking respondents how many people they know in the general population, given some definition of what it means to know another person. Since previous studies have found these responses to be inaccurate (Killworth2003; McCormick et al., 2010), researchers often estimate the degrees instead by asking other questions. Therefore, we can decompose the estimation problem into two steps: first estimating respondents’ degrees, then using these degree estimates $\hat{D}_{i}$ with the responses $Y_{iH}$ to estimate the prevalence $r$.
The NSUM approach is based on the idea that given the response $Y_{iH}$ and degree $D_{i}$ for one person, a rough estimate of the hard-to-reach group prevalence is $Y_{iH}/D_{i}$ (Bernard et al., 1991). To obtain a better estimate, we can pool the responses and degrees from a larger sample of people and model the responses $Y_{iH}$ as binomially distributed; henceforth we refer to this early NSUM model as the binomial model (Killworth et al., 1998). This information can be pooled in two basic ways, either the ratio of average response to average degree or the average ratio of response to degree (hereafter, we refer to the ratio of averages of a quantity as R and the average of ratios as A):
@@
A systematic literature review by Ocagli2021 of all PubMed papers from the original NSUM paper through 2020 that use NSUM ultimately includes 35 studies. Examining these papers ourselves, we found that five of them focus on methods development and another two only estimate network size, not population size. Of the remaining 28 studies, all appear to use the $*{}^{*}$ estimator, sometimes with adjustment factors; one compares against generalized NSUM, another compares against the AA estimator, and some are not explicit about the specific estimator or procedure. None of these studies use or consider the RA or AR estimators. The most commonly used of these four estimators therefore seems to be the $*{}^{*}$ estimator, first proposed to our knowledge by Killworth et al. (1998). Throughout this paper, we write $*{}^{*}$ as a reminder that this is the current default basic NSUM estimator.
The goal of the present study is to systematically evaluate the comparative performance of these estimators. An earlier study by Killworth et al. (1998) compares the $*{}^{*}$ and RA estimators, and a more recent paper by Habecker et al. (2015) promotes the use of the AA estimator, which they call the mean-of-sums (MoS) estimator. Habecker et al. (2015) suggest potential motivations for using the AA estimator and demonstrate on an example data set that the $*{}^{*}$ and AA estimators (as well as other modifications they propose) result in very different estimates, but they do not provide a clear theoretical or empirical evaluation of which estimator seems to be less biased. Laga et al. (2021) refer to the RA estimator as the plug-in MLE, or PIMLE. We are not aware of any literature considering or proposing the use of the AR estimator, but we evaluate it in the present study along with the other three estimators.
When the binomial model is true, both the A and R prevalence estimators with fixed or known degrees are unbiased, but the latter has a smaller variance; we provide a proof in the online supplement using reasoning about harmonic and arithmetic means. Killworth et al. (1998) compute and empirically evaluate the $*{}^{*}$ and RA on a data set and state that the latter has an unacceptably high variance, but their theoretical analysis assumes the degrees are known; thus their theoretical analysis concerns only the (one-step) A and R prevalence estimators, not the (two-step) $*{}^{*}$ and RA estimators. The degree estimation step in the $*{}^{*}$ and RA estimators not only involves additional uses of the conditional proportion assumption to estimate the degree, but also accounts for the distribution of degrees rather than conditioning on them. Perhaps this is the reason that surveys tend to use the $*{}^{*}$ if they use one of the simple estimators, and that researchers tend to start from the $*{}^{*}$ when they develop methods to extend the NSUM approach and relax modeling assumptions.
@@ 3 A framework for studying estimator behavior under barrier effects
Killworth et al. (1998) assume a simple random sample without replacement, while Feehan2015Thesis and Habecker et al. (2015) propose a way to extend this to general probability survey designs. For simplicity and to stay consistent with the original estimators, in this study we assume simple random sampling without replacement.
@@ 4 Analytical results on estimator bias and variance
Note that when the three link probabilities are equal, corresponding to the binomial model, the $*{}^{*}$ estimator does not have smaller variance as suggested by Killworth et al. (1998): both estimators have the same first-order variance. We believe this is the first time this result has been shown for these estimators. In this case, the first-order approximations of both estimators’ expectations equal the true prevalence; in the language of Feehan and Salganik (2016) and other literature, both the $*{}^{*}$ and RA estimators are essentially unbiased.

### 17 Jan 2024 v3

@@ 1 Introduction
Surveys are a standard approach to estimating the size of subpopulations, or groups of people with a particular trait. Many key subpopulations of interest, however, are hard to reach***We note that researchers and community partners have raised concerns about the term “hard-to-reach”; we use this term here to encompass communities that match one or more of the above conditions, and we echo the message of Routen et al. (2022) that the onus of reaching such populations is on researchers and funders, not on the populations themselves. with standard surveys for a number of possible reasons: (1) it may be hard to get a list of the members of this subpopulation or they may be hard to contact, as in the case of the homeless subpopulation; (2) it may be hard to accurately determine their membership in the subpopulation because their group status is stigmatized, as in the case of heavy drug users; and/or (3) the subpopulation may be rare in the general population (Bernard et al., 1991; Killworth et al., 1998a). The limitations of standard survey methodology in reaching these groups has motivated the development of better adapted methods.
@@
The other strategy, the network scale-up method (NSUM), is to conduct a traditional survey with a representative sample from the general or frame population and ask respondents how many people they know in the hard-to-reach populations of interest, then use information about their personal network sizes (degrees) to scale up their data into an estimate for the general prevalence of those populations (Bernard et al., 1991, 2010). Researchers could ask respondents how many people they know (their degree) and how many people they know in the hard-to-reach group. However, people often cannot provide an accurate estimate for their degree (Killworth et al., 2003; McCormick et al., 2010), leading to inaccurate hard-to-reach population prevalence estimates. Therefore, respondents’ degrees are often estimated by asking other questions that are used to estimate the degree. In the probe group method, respondents are asked how many people they know in various groups of known size that collectively are thought to be representative of the general population (Killworth et al., 1998a; McCarty et al., 2001; McCormick et al., 2010). The responses to questions of the form “How many people do you know with X trait?” are known as aggregated relational data or ARD (McCormick and Zheng, 2015). This approach does not require knowing whether the respondents themselves are in the hard-to-reach populations.
In this paper we focus on the NSUM strategy. A recent literature review by Ocagli et al. (2021) suggests the majority of studies use one of the simplest estimators. However, we find that when a key model assumption is violated by the presence of barrier effects, which may often occur in practice, the mean squared error (MSE) may actually tend to be much smaller for another simple estimator that is equally easy to implement but much less commonly used. We demonstrate this through theoretical derivations as well as simulated surveys on a real network data set.
@@ 2 NSUM estimators
The target of estimation is the prevalence $r$ of the hard-to-reach group $H$ in the general population, which equals the ratio of the group size $N_{H}$ to the size $N$ of the general or frame population. For each respondent in a sample of size $n$, let $Y_{iH}$ represent respondent $i$’s response to the ARD question, “How many people do you know in $H$?” Let $D_{i}$ represent the degree of respondent $i$. In principle, we could determine $D_{i}$ by directly asking respondents how many people they know in the general population, given some definition of what it means to know another person. Since previous studies have found these responses to be inaccurate (Killworth et al., 2003; McCormick et al., 2010), researchers often estimate the degrees instead by asking other questions. Therefore, we can decompose the estimation problem into two steps: first estimating respondents’ degrees, then using these degree estimates $\hat{D}_{i}$ with the responses $Y_{iH}$ to estimate the prevalence $r$.
The NSUM approach is based on the idea that given the response $Y_{iH}$ and degree $D_{i}$ for one person, a rough estimate of the hard-to-reach group prevalence is $Y_{iH}/D_{i}$ (Bernard et al., 1991). To obtain a better estimate, we can pool the responses and degrees from a larger sample of people and model the responses $Y_{iH}$ as binomially distributed; henceforth we refer to this early NSUM model as the binomial model (Killworth et al., 1998a). This information can be pooled in two basic ways, either the ratio of average response to average degree or the average ratio of response to degree (hereafter, we refer to the ratio of averages of a quantity as R and the average of ratios as A):
@@
A systematic literature review by Ocagli et al. (2021) of all PubMed papers from the original NSUM paper through 2020 that use NSUM ultimately includes 35 studies. Examining these papers ourselves, we found that five of them focus on methods development and another two only estimate network size, not population size. Of the remaining 28 studies, all appear to use the $*{}^{*}$ estimator, sometimes with adjustment factors; one compares against generalized NSUM, another compares against the AA estimator, and some are not explicit about the specific estimator or procedure. None of these studies use or consider the RA or AR estimators. The most commonly used of these four estimators therefore seems to be the $*{}^{*}$ estimator, first proposed to our knowledge by Killworth et al. (1998b). Throughout this paper, we write $*{}^{*}$ as a reminder that this is the current default basic NSUM estimator.
The goal of the present study is to systematically evaluate the comparative performance of these estimators. An earlier study by Killworth et al. (1998a) compares the $*{}^{*}$ and RA estimators, and a more recent paper by Habecker et al. (2015) promotes the use of the AA estimator, which they call the mean-of-sums (MoS) estimator. Habecker et al. (2015) suggest potential motivations for using the AA estimator and demonstrate on an example data set that the $*{}^{*}$ and AA estimators (as well as other modifications they propose) result in very different estimates, but they do not provide a clear theoretical or empirical evaluation of which estimator seems to be less biased. Laga et al. (2021) refer to the RA estimator as the plug-in MLE, or PIMLE. We are not aware of any literature considering or proposing the use of the AR estimator, but we evaluate it in the present study along with the other three estimators.
When the binomial model is true, both the A and R prevalence estimators with fixed or known degrees are unbiased, but the latter has a smaller variance; we provide a proof in the online supplement using reasoning about harmonic and arithmetic means. Killworth et al. (1998a) compute and empirically evaluate the $*{}^{*}$ and RA on a data set and state that the latter has an unacceptably high variance, but their theoretical analysis assumes the degrees are known; thus their theoretical analysis concerns only the (one-step) A and R prevalence estimators, not the (two-step) $*{}^{*}$ and RA estimators. The degree estimation step in the $*{}^{*}$ and RA estimators not only involves additional uses of the conditional proportion assumption to estimate the degree, but also accounts for the distribution of degrees rather than conditioning on them. Perhaps this is the reason that surveys tend to use the $*{}^{*}$ if they use one of the simple estimators, and that researchers tend to start from the $*{}^{*}$ when they develop methods to extend the NSUM approach and relax modeling assumptions.
@@ 3 A framework for studying estimator behavior under barrier effects
Killworth et al. (1998a) assume a simple random sample without replacement, while Feehan (2015) and Habecker et al. (2015) propose a way to extend this to general probability survey designs. For simplicity and to stay consistent with the original estimators, in this study we assume simple random sampling without replacement.
@@ 4 Analytical results on estimator bias and variance
Note that when the three link probabilities are equal, corresponding to the binomial model, the $*{}^{*}$ estimator does not have smaller variance as suggested by Killworth et al. (1998a): both estimators have the same first-order variance. We believe this is the first time this result has been shown for these estimators. In this case, the first-order approximations of both estimators’ expectations equal the true prevalence; in the language of Feehan and Salganik (2016) and other literature, both the $*{}^{*}$ and RA estimators are essentially unbiased.